# On the Complexity of Determinations

### 30 Mar 2026 v1

@@ C.1 Conservation of sequential depth (Theorem 3)
The constrained generation task is offline: the constraint chain is fully given, so there is one complete history with no extensions. This rules out branch-specific conditioning—the strategy cannot make different choices for different futures, because there are no branches. The following remark shows why this matters.
Remark 4 (Branch-specific commitments can bypass conservation).
In the online setting, a basis that allows branch-specific commitments—whose effects depend on which future history materializes—can bypass the conservation bound. In the three-valued consensus of Example 1 (an online specification), $\mathsf{Depth}_{\Phi}(\mathsf{Spec})=2$ (under the atomic basis), but a single non-pointwise commitment can choose differently per branch (keep $b$ at $H_{1}$, $c$ at $H_{2}$, $c$ at $H_{3}$), resolving the specification in one layer with $O(1)$ circuit depth.
The offline product-space setting of the constrained generation task rules out this escape: all commitments are applied after the history is fixed, and the coordinate basis is commutative, so no branch-specific conditioning is possible.
C.2 Tightness of the conservation law and the three-way tradeoff
@@ Appendix D Applications to Classical Problems
This section demonstrates the breadth of the determination framework by applying it to problems across several domains: stable matching, extensive-form games, chain-of-thought reasoning, and distributed round complexity (BSP and LOCAL). In each case, the framework either recovers a known result with a new explanation or reveals structure that existing models do not capture. Determination depth arises for two reasons across these examples: enablement, where each commitment creates the sub-problem for the next layer (constrained generation, extensive-form games); and interference, where commitments that are individually valid conflict when applied in the same layer (stable matching under the rotation basis). Each example connects to known results in its domain: stable matching recovers Garg’s parallel algorithm [GAR20] and witnesses the orthogonality of determination and computational depth (finding a stable matching is in P, yet determination depth can be arbitrarily large); extensive-form games measure the game’s strategic depth—the moves where a player must break a tie among equally-optimal options—connecting to Selten’s trembling-hand refinement [SEL75]; chain-of-thought reasoning decomposes CoT length into determination depth, computational depth, and architectural overhead, explaining why longer chains sometimes degrade [SMA+25, SZW+25]; and BSP/LOCAL expose hidden modeling assumptions (Ameloot et al.’s non-obliviousness [ANd13]) that silently collapse determination layers.
@@ D.1.2 Determination depth equals rotation poset height
By the Irving–Leather realization theorem [GI89], every finite poset arises as a rotation poset. Since determination depth equals poset height (Proposition 6), stable matching realizes every possible determination depth: for every $k$, there exists an instance with determination depth exactly $k$. In particular, finding a stable matching is in P, yet the determination depth can be arbitrarily large—computational cost is polynomial regardless of determination depth.
@@
Whether the conservation law yields new lower bounds on parallel stable matching algorithms is an open question.
@@ D.2 Strategic Depth in Extensive-Form Games
For a two-player extensive-form game with perfect information, the determination depth from player 1’s perspective (under the SPE-restricted specification) equals the maximum number of subgame-non-trivial player-1 nodes on any root-to-leaf path.
@@
Lower bound. Consider two consecutive subgame-non-trivial player-1 nodes $u$ and $v$ on a root-to-leaf path, with $v$ deeper than $u$. Node $v$ has SPE-consistent children $c_{1},c_{2}$, so there exist SPEs $\sigma_{1},\sigma_{2}$ of the full game inducing plays through $c_{1}$ and $c_{2}$ respectively. Between $u$ and $v$, player-2 moves and trivial player-1 moves occur as environment events. The SPEs $\sigma_{1}$ and $\sigma_{2}$ may prescribe different player-2 moves at nodes between $u$ and $v$, so the admissible set at vv—which plays remain SPE-consistent—depends on which environment events materialize after the commitment at $u$. Under one sequence of player-2 moves, only plays through $c_{1}$ may remain admissible at $v$; under another, only plays through $c_{2}$. Since these environment events have not occurred when the commitment at $u$ is made, no commitment at $u$ can validly determine the choice at $v$: any fixed choice risks selecting a child whose plays are inadmissible under some SPE-consistent continuation. The choice at $v$ is therefore a forced dependency on the history up to $v$. Chaining these forced dependencies across all non-trivial nodes on the path gives a dependency chain of length equal to their count (Proposition 4). ∎
The result decomposes game-tree depth into two components: strategic depth (the non-trivial nodes, where the player must break a tie among equally-optimal moves) and forced depth (the trivial nodes, where the optimal move is uniquely determined and no choice is involved). This decomposition is distinct from classical game-tree depth and alternating Turing machine quantifier depth, both of which count all player-1 moves—including forced ones that involve no relational choice. The distinction is analogous to the BSP diagnostic (Section D.4): just as BSP overcharges by counting communication rounds that resolve no ambiguity, game-tree depth overcharges by counting moves that present no relational choice. The analysis is restricted to finite perfect-information games (Zermelo’s setting); extending to imperfect information requires handling information sets and is left open.
@@
Beyond extensive-form games, the determination framework applies to mechanism design. A mechanism (e.g., an auction format) can be viewed as enriching the commitment basis: a direct-revelation mechanism allows a player to submit a full strategy as a single commitment, collapsing multiple sequential choices into one layer. A dominant-strategy mechanism (e.g., a second-price auction [VIC61]) makes every node subgame-trivial—each player’s optimal action is independent of others’—collapsing strategic depth to 0. An indirect mechanism (e.g., an ascending auction [MIL00]) trades strategic depth for communication: more rounds of bidding, but less information per round. This echoes the depth–width–communication tradeoff (Theorem 2): formalizing the precise relationship—whether indirect mechanisms obey an analogous conservation law trading determination depth against per-round communication and outcome-set width—is an open problem.
@@ D.3 Decomposing Chain-of-Thought Length
We use the determination framework to decompose CoT length into formally independent components. The decomposition is not a theorem about any specific transformer architecture; it is a semantic lower bound that applies to any autoregressive procedure in which each token is an irrevocable commitment and the task is relational (multiple valid continuations exist). The constrained generation task (Section 3) serves as the witness because it isolates determination depth from computational depth: each commitment is a constant-time table lookup, so the sequential cost is entirely semantic. The total CoT length $T$ is lower-bounded by three formally independent quantities: (i) determination depth dd—non-commuting commitment layers, witnessed by the constrained generation task (Section 3); (ii) computational depth cc—inherent per-step sequentiality from bounded transformer depth (TC0-like [FZG+23]), witnessed by graph connectivity [MAA+25]; and (iii) architectural overhead—tokens spent on information management (context summarization, backtracking, re-derivation), reducible by changing the architecture without changing the task.
@@ Appendix F Related Work
This section positions determination depth among existing notions of staged computation and semantic dependency. Closest antecedents treat staging either as an operational resource (rounds, adaptivity, oracle calls) or as a syntactic discipline (stratification, fixpoint nesting). Our contribution is to identify staging as a semantic invariant of resolution—defined prior to any particular model or language.
@@
Depth as a complexity measure has a long history in models of parallel computation. Circuit complexity distinguishes size from depth, isolating irreducible sequential structure even when unbounded parallelism is available [PIP80, BAR90]. Circuit depth measures the inherent sequentiality of evaluating a function: data dependencies force some gates to wait for others. Determination depth measures the inherent sequentiality of choosing an outcome from a relation: commitment dependencies force some decisions to wait for others. The two are orthogonal: a function has determination depth 0 regardless of its circuit depth, while a relation can have high determination depth even when every individual commitment is a constant-depth circuit (Section 3). This orthogonality is witnessed in the offline setting by the constrained generation task (Section 3); in the online setting, forward validity constraints add a further source of determination depth that circuit models do not capture at all. PRAM and BSP models similarly separate local computation from global synchronization, charging depth to barriers or rounds [BC82, VAL90]. In the online (distributed) setting, these models can both undercharge (when a round boundary hides an irrevocable commitment that the model treats as primitive) and overcharge (when communication structure is conflated with semantic resolution); Section D.4 develops this diagnostic in detail. More broadly, classical hierarchy theorems establish strict separations based on time, space, or alternation depth [SIP83]; our determination depth hierarchies are orthogonal, arising from semantic dependency rather than resource bounds.
@@
The layered normal form (Section A) resembles the Foata normal form of Mazurkiewicz trace theory [MAZ77]: a canonical factorization of a word in a partially commutative monoid into maximal independent steps. The resemblance is structural but the theories diverge in three ways that produce qualitatively different phenomena. First, in classical trace theory the independence relation is fixed: two letters either commute or they do not, regardless of context. In determination theory, commutation is dynamic: whether two commitments commute depends on the current refined specification, and applying one commitment can create or destroy independence among the remaining ones. This dynamic commutation arises for different reasons in the two settings: in the online setting, forward validity constraints (a commitment that is safe alone may become unsafe after another commitment narrows the admissible set at some extension) create and destroy independence; in the offline setting, enablement (committing to one position determines which choices remain feasible at the next) produces the same effect. Second, this state-dependence makes the analogue of the Foata normal form non-unique—different layering choices lead to different refined specifications and potentially different minimum determination depths—and computing the minimum height becomes NP-hard (Proposition 9), in contrast to the polynomial-time computability of static Foata height. Third, the main results of this paper—the exponential depth–width separation (offline), the oracle characterization (online), and the conservation law (both settings)—have no analogues in classical trace theory. They arise from the semantic content of commitments (feasibility, resolution, forward validity) rather than from the algebraic structure of partial commutation alone. In short, trace theory provides the algebraic skeleton; the semantic content of determination fills it with phenomena that the skeleton alone cannot express.
@@
The use of a disclosure oracle to characterize depth parallels classical distinctions between adaptive and non-adaptive computation. Decision tree complexity, oracle Turing machines, and communication complexity all exhibit hierarchies based on the number of adaptive rounds [KL80, PS84, NIS91]. Our oracle characterization (Proposition 1) is specific to the online setting under a commutative basis: oracle invocations witness irreducible semantic dependency—points at which the strategy must observe the evolving history before committing further—rather than query adaptivity or information flow. Local computation between oracle calls is unrestricted, emphasizing that determination depth cannot be collapsed by parallelism or control flow alone. In the offline setting, a single oracle call reveals the complete input and the oracle count collapses to one; the sequential cost reappears as computational depth within layers (the conservation law, Theorem 3). In this sense, determination depth and adaptivity depth both measure staged dependency, but the dependency has a different character. An adaptive query reveals information about a fixed, predetermined answer; a determination commitment creates the answer by irrevocably excluding alternatives. Adaptivity depth measures how many times a strategy must look; determination depth measures how many times it must choose. The two are orthogonal (Observation 3, Section 3): pointer chasing at sparsity $s=1$ has high adaptivity depth but determination depth zero (the answer is unique); constrained generation in the offline setting has determination depth $k$ but adaptivity depth zero (the entire input is given).
@@
Layered semantics appear in database theory through stratified negation, well-founded semantics, and alternating fixpoints [GRS91]; in descriptive complexity, fixpoint nesting and alternation depth classify expressive power [IMM99]. These layerings are properties of a program under a chosen semantics: a stratified Datalog program with $k$ strata requires $k$ non-commuting sealing commitments (one per stratum), but applying well-founded or stable semantics to the same program yields a different commitment structure and potentially a different determination depth. Determination depth, by contrast, is a property of the specification (the mapping from histories to admissible outcome sets), independent of any particular logical formalism or semantics used to define it. Two logic programs that define the same specification have the same determination depth, even if they differ in strata count, fixpoint nesting, or use of negation. Conversely, applying different negation semantics to the same program can yield different specifications and hence different determination depths—a distinction that strata counts alone cannot express.
@@
The leaf language framework of Bovet, Crescenzi, and Silvestri [BCS92] characterizes complexity classes by the language-theoretic complexity of the string of outcomes at the leaves of a nondeterministic computation tree: NP corresponds to testing membership in $\{0,1\}^{*}1\{0,1\}^{*}$, PSPACE to a regular language recognizable with $O(\log n)$ memory, and so on. Both frameworks use the structure of a tree of possibilities to classify problems, but the objects and questions differ. Leaf languages classify the acceptance condition applied to a fixed nondeterministic computation; determination depth classifies the commitment structure required to resolve a relational specification. In particular, leaf languages operate on a single computation tree whose branching is fixed by the machine, while determination depth measures layered commitment across histories whose extensions are chosen by an adversarial environment—a distinction that is specific to the online setting. The PH characterization (Theorem 4) recovers the same hierarchy that leaf languages capture, but through alternation of commitment and environment moves in the online game rather than through the complexity of a leaf-string acceptance condition. In the offline setting, the adversarial environment disappears and the exponential separation (Theorem 1) provides a different route to determination depth hierarchies, via commitment dependency rather than alternation. More broadly, the leaf language program provided an elegant classification of existing complexity classes but did not yield new separations or lower bounds. Our framework is aimed at a different target: not reclassifying known classes, but identifying an axis of complexity orthogonal to computation (Section 3), producing new tradeoffs (Theorems 12), and diagnosing existing models (Sections $D.4$).
@@
A growing body of work studies the power and limitations of chain-of-thought reasoning in transformers [FZG+23, LLZ+24, MAA+25]. Feng, Zhang, Gu, Ye, He, and Wang [FZG+23] and Li and Liu [LLZ+24] show that CoT enables bounded-depth transformers (TC0) to solve problems requiring greater computational depth. Mirtaheri, Anil, Agarwal, and Neyshabur [MAA+25] prove an exponential separation between sequential and parallel scaling for graph connectivity. These results concern computational depth—the inherent sequentiality of evaluating a function. Our exponential separation (Section 3) establishes an orthogonal source of sequential advantage: determination depth, arising from non-commuting commitments in relational specifications. The separation is an offline result (the full constraint chain is given); the decomposition of Appendix D.3 extends to the online autoregressive setting, unifying computational depth, determination depth, and architectural overhead into a single framework. Complementary empirical work identifies failure modes of extended reasoning [SMA+25, SZW+25] and studies optimal compute allocation [SLX+24]; our decomposition provides a formal account of why these phenomena arise.
@@
Across these domains, notions of depth have appeared as proxies for sequentiality, staging, or adaptivity. Our contribution is to identify determination depth and determination cost as semantic complexity measures—arising from forward validity constraints in the online setting and from commitment dependency in the offline setting—that are orthogonal to computational complexity. Determination depth measures the minimum sequential layers of irrevocable commitment; determination cost measures the total number of commitments; the exponential separation (Theorem 1) shows that reducing one without increasing the other is impossible.

### 01 Apr 2026 v2

@@ Appendix D Applications to Classical Problems
This section demonstrates the breadth of the determination framework by applying it to problems across several domains: stable matching, extensive-form games, chain-of-thought reasoning, and distributed round complexity (BSP and LOCAL). In each case, the framework either recovers a known result with a new explanation or reveals structure that existing models do not capture. Determination depth arises for two reasons across these examples: commitment dependency, where each commitment creates the sub-problem for the next layer (constrained generation, extensive-form games); and forward validity, where commitments that are individually valid conflict when applied in the same layer (stable matching under the rotation basis). Each example connects to known results in its domain: stable matching recovers Garg’s parallel algorithm [GAR20] and witnesses the universality of determination depth (every finite depth is realized by some instance); extensive-form games measure the game’s strategic depth—the moves where a player must break a tie among equally-optimal options—connecting to Selten’s trembling-hand refinement [SEL75]; chain-of-thought reasoning decomposes CoT length into determination depth, computational depth, and architectural overhead, explaining why longer chains sometimes degrade [SMA+25, SZW+25]; and BSP/LOCAL expose hidden modeling assumptions (Ameloot et al.’s non-obliviousness [ANd13]) that silently collapse determination layers.
@@ D.1.2 Determination depth equals rotation poset height
By the Irving–Leather realization theorem [GI89], every finite poset arises as a rotation poset. Since determination depth equals poset height (Proposition 5), stable matching realizes every possible determination depth: for every $k$, there exists an instance with determination depth exactly $k$.
@@ D.2 Strategic Depth in Extensive-Form Games
For a two-player extensive-form game with perfect information, the determination depth from player 1’s perspective (under the SPE-restricted specification) is at most the maximum number of subgame-non-trivial player-1 nodes on any root-to-leaf path. The bound is tight whenever, between each consecutive pair of such nodes, player 2 has at least one subgame-non-trivial node.
@@
Lower bound (under the tightness condition). Consider two consecutive subgame-non-trivial player-1 nodes $u$ and $v$ on a root-to-leaf path, with $v$ deeper than $u$. Node $v$ has SPE-consistent children $c_{1},c_{2}$, so there exist SPEs $\sigma_{1},\sigma_{2}$ of the full game inducing plays through $c_{1}$ and $c_{2}$ respectively. Between $u$ and $v$, player-2 moves and trivial player-1 moves occur as environment events. The SPEs $\sigma_{1}$ and $\sigma_{2}$ may prescribe different player-2 moves at nodes between $u$ and $v$, so the admissible set at vv—which plays remain SPE-consistent—depends on which environment events materialize after the commitment at $u$. Under one sequence of player-2 moves, only plays through $c_{1}$ may remain admissible at $v$; under another, only plays through $c_{2}$. Since these environment events have not occurred when the commitment at $u$ is made, no commitment at $u$ can validly determine the choice at $v$: any fixed choice risks selecting a child whose plays are inadmissible under some SPE-consistent continuation. The choice at $v$ is therefore a forced dependency on the history up to $v$. Chaining these forced dependencies across all non-trivial nodes on the path gives a dependency chain of length equal to their count (Proposition 3). ∎
The result decomposes game-tree depth into two components: strategic depth (the non-trivial nodes, where the player must break a tie among equally-optimal moves) and forced depth (the trivial nodes, where the optimal move is uniquely determined and no choice is involved). This decomposition is distinct from classical game-tree depth and alternating Turing machine quantifier depth, both of which count all player-1 moves—including forced ones that involve no relational choice. The distinction is analogous to the BSP diagnostic (Section D.4): just as BSP overcharges by counting communication rounds that resolve no ambiguity, game-tree depth overcharges by counting moves that present no relational choice. The analysis is restricted to finite perfect-information games (Zermelo’s setting [ZER13]); extending to imperfect information requires handling information sets and is left open.
@@
Beyond extensive-form games, the determination framework applies to mechanism design. A mechanism (e.g., an auction format) can be viewed as enriching the commitment basis: a direct-revelation mechanism allows a player to submit a full strategy as a single commitment, collapsing multiple sequential choices into one layer. A dominant-strategy mechanism (e.g., a second-price auction [VIC61]) makes every node subgame-trivial—each player’s optimal action is independent of others’—collapsing strategic depth to 0. An indirect mechanism (e.g., an ascending auction [MIL00]) trades strategic depth for communication: more rounds of bidding, but less information per round. Whether indirect mechanisms obey a conservation law analogous to Theorem 3—trading determination depth against per-round communication and outcome-set width—is an open problem.
@@ D.3 Decomposing Chain-of-Thought Length
We use the determination framework to decompose CoT length into formally independent components. The decomposition is not a theorem about any specific transformer architecture; it is a semantic lower bound that applies to any autoregressive procedure in which each token is an irrevocable commitment and the task is relational (multiple valid continuations exist). The constrained generation task (Section 3) serves as the witness because it isolates determination depth from computational depth: each commitment is a constant-time table lookup, so the sequential cost is entirely semantic. The total CoT length $T$ is lower-bounded by three independent quantities: (i) determination depth dd—non-commuting commitment layers, witnessed by the constrained generation task (Section 3); (ii) computational depth cc—inherent per-step sequentiality from bounded transformer depth (TC0-like [FZG+23]), witnessed by graph connectivity [MAA+25]; and (iii) architectural overhead—tokens spent on information management (context summarization, backtracking, re-derivation), reducible by changing the architecture without changing the task.
@@ Appendix F Related Work
This section positions determination depth relative to existing complexity measures that involve sequential staging. These measures treat staging either as an operational resource (rounds, adaptivity, oracle calls) or as a syntactic discipline (stratification, fixpoint nesting); determination depth differs in that it measures the semantic cost of irrevocable commitment, independent of any particular model or language.
@@
Depth as a complexity measure has a long history in models of parallel computation. Circuit complexity distinguishes size from depth, isolating irreducible sequential structure even when unbounded parallelism is available [PIP80, BAR90]. Circuit depth measures the inherent sequentiality of evaluating a function: data dependencies force some gates to wait for others. Determination depth measures the inherent sequentiality of choosing an outcome from a relation: commitment dependencies force some decisions to wait for others. The two are distinct: a function has determination depth 0 regardless of its circuit depth, while a relation can have high determination depth even when every individual commitment is a constant-depth circuit (Section 3). In the online setting, forward validity constraints add a further source of determination depth that circuit models do not capture at all. PRAM and BSP models similarly separate local computation from global synchronization, charging depth to barriers or rounds [BC82, VAL90]. In the online (distributed) setting, these models can both undercharge (when a round boundary hides an irrevocable commitment that the model treats as primitive) and overcharge (when communication structure is conflated with semantic resolution); Section D.4 develops this diagnostic in detail. More broadly, classical hierarchy theorems establish strict separations based on time, space, or alternation depth [SIP83]; our hierarchies arise from semantic dependency rather than resource bounds.
@@
The layered normal form (Section A) resembles the Foata normal form of Mazurkiewicz trace theory [MAZ77]: a canonical factorization of a word in a partially commutative monoid into maximal independent steps. The resemblance is structural but the theories diverge in three ways that produce qualitatively different phenomena. First, in classical trace theory the independence relation is fixed: two letters either commute or they do not, regardless of context. In determination theory, commutation is dynamic: whether two commitments commute depends on the current refined specification, and applying one commitment can create or destroy independence among the remaining ones. This dynamic commutation arises for different reasons in the two settings: in the online setting, forward validity constraints (a commitment that is safe alone may become unsafe after another commitment narrows the admissible set at some extension) create and destroy independence; in the offline setting, commitment dependency (committing to one position determines which choices remain feasible at the next) produces the same effect. Second, this state-dependence makes the analogue of the Foata normal form non-unique—different layering choices lead to different refined specifications and potentially different minimum determination depths—and computing the minimum height becomes NP-hard (Proposition 8), in contrast to the polynomial-time computability of static Foata height. Third, the main results of this paper—the exponential depth–width separation (offline), the oracle characterization (online), and the conservation law (both settings)—have no analogues in classical trace theory. They arise from the semantic content of commitments (feasibility, resolution, forward validity) rather than from the algebraic structure of partial commutation alone. In short, trace theory provides the algebraic skeleton; the semantic content of determination fills it with phenomena that the skeleton alone cannot express.
@@
The use of a disclosure oracle to characterize depth parallels classical distinctions between adaptive and non-adaptive computation. Decision tree complexity, oracle Turing machines, and communication complexity all exhibit hierarchies based on the number of adaptive rounds [KL80, PS84, NIS91]. Our oracle characterization (Proposition 1) is specific to the online setting under a commutative basis: oracle invocations witness irreducible semantic dependency—points at which the strategy must observe the evolving history before committing further—rather than query adaptivity or information flow. Local computation between oracle calls is unrestricted, emphasizing that determination depth cannot be collapsed by parallelism or control flow alone. In the offline setting, a single oracle call reveals the complete input and the oracle count collapses to one; the sequential cost reappears as computational depth within layers (the conservation law, Theorem 3). In this sense, determination depth and adaptivity depth both measure staged dependency, but the dependency has a different character. An adaptive query reveals information about a fixed, predetermined answer; a determination commitment creates the answer by irrevocably excluding alternatives. Adaptivity depth measures how many times a strategy must look; determination depth measures how many times it must choose. The two are independent (Observation 3, Section 3): pointer chasing at sparsity $s=1$ has high adaptivity depth but determination depth zero (the answer is unique); constrained generation in the offline setting has determination depth $k$ but adaptivity depth zero (the entire input is given).
@@
Layered semantics appear in database theory through stratified negation, well-founded semantics, and alternating fixpoints [GRS91]; in descriptive complexity, fixpoint nesting and alternation depth classify expressive power [IMM99]. These layerings are properties of a program under a chosen semantics. Determination depth, by contrast, is a property of the specification (the mapping from histories to admissible outcome sets), independent of any particular logical formalism or semantics: two programs that define the same specification have the same determination depth, even if they differ in strata count, fixpoint nesting, or use of negation. The distinction matters because different negation semantics applied to the same program can yield different specifications and hence different determination depths. Each semantics induces a different commitment basis—stratified semantics requires a linear chain of sealing commitments, well-founded semantics an alternating sequence, stable semantics a shallow branching among incompatible determinations—a distinction strata counts alone cannot express.
@@
The leaf language framework of Bovet, Crescenzi, and Silvestri [BCS92] characterizes complexity classes by the language-theoretic complexity of the string of outcomes at the leaves of a nondeterministic computation tree: NP corresponds to testing membership in $\{0,1\}^{*}1\{0,1\}^{*}$, PSPACE to a regular language recognizable with $O(\log n)$ memory, and so on. Both frameworks use the structure of a tree of possibilities to classify problems, but the objects and questions differ. Leaf languages classify the acceptance condition applied to a fixed nondeterministic computation; determination depth classifies the commitment structure required to resolve a relational specification. In particular, leaf languages operate on a single computation tree whose branching is fixed by the machine, while determination depth measures layered commitment across histories whose extensions are chosen by an adversarial environment—a distinction that is specific to the online setting. The PH characterization (Theorem 4) recovers the same hierarchy that leaf languages capture, but through alternation of commitment and environment moves in the online game rather than through the complexity of a leaf-string acceptance condition. In the offline setting, the adversarial environment disappears and the exponential separation (Theorem 1) provides a different route to determination depth hierarchies, via commitment dependency rather than alternation. More broadly, the leaf language program provided an elegant classification of existing complexity classes but did not yield new separations or lower bounds. Our framework is aimed at a different target: not reclassifying known classes, but identifying an axis of complexity distinct from computation (Section 3), producing new tradeoffs (Theorems 12), and diagnosing existing models (Sections $D.4$).
@@
A growing body of work studies the power and limitations of chain-of-thought reasoning in transformers [FZG+23, LLZ+24, MAA+25]. Feng, Zhang, Gu, Ye, He, and Wang [FZG+23] and Li and Liu [LLZ+24] show that CoT enables bounded-depth transformers (TC0) to solve problems requiring greater computational depth. Mirtaheri, Anil, Agarwal, and Neyshabur [MAA+25] prove an exponential separation between sequential and parallel scaling for graph connectivity. These results concern computational depth—the inherent sequentiality of evaluating a function. Our exponential separation (Section 3) establishes a distinct source of sequential advantage: determination depth, arising from non-commuting commitments in relational specifications. The separation is an offline result (the full constraint chain is given); the decomposition of Appendix D.3 extends to the online autoregressive setting, unifying computational depth, determination depth, and architectural overhead into a single framework. Complementary empirical work identifies failure modes of extended reasoning [SMA+25, SZW+25] and studies optimal compute allocation [SLX+24]; our decomposition provides a formal account of why these phenomena arise.
@@
Across these domains, notions of depth have appeared as proxies for sequentiality, staging, or adaptivity. Our contribution is to identify determination depth and determination cost as semantic complexity measures—arising from forward validity constraints in the online setting and from commitment dependency in the offline setting—that are distinct from computational complexity. Determination depth measures the minimum sequential layers of irrevocable commitment; determination cost measures the total number of commitments; the exponential separation (Theorem 1) shows that reducing one without increasing the other is impossible.